# Lattice isomorphic Banach lattices of polynomials

### v1

@@ 2 Díaz-Dineen problem for regular polynomials
is positive, symmetric (because $\overline{T_{P}}$ is symmetric by the assumption and Lemma 2.2) and satisfies $\overline{\varphi_{G}}(P)(y)=A(y,\dots,y)$. Thus $\overline{\varphi_{G}}(P)\geq 0$ for every positive $n$-homogeneous polynomial $P\colon E\to G$. Therefore, we may consider the positive linear operator $\overline{\varphi_{G}}$ from $\mathcal{P}^{r}(^{n}E;G^{\ast})$ into $\mathcal{P}^{r}(^{n}F;G^{\ast})$. Analogously, since $\varphi^{-1}\colon F^{\ast}\to E^{\ast}$ is a positive operator (see [1, Theorem 2.15]), we get that $\overline{\varphi_{G}^{-1}}\colon\mathcal{P}^{r}(^{n}F;G^{\ast})\to\mathcal{P}^{r}(^{n}E;G^{\ast})$ is also a positive operator.
Finally, since $E^{*}$ and $F^{*}$ have order continuous norm, we get by the same argument used in the proof of [14, Theorem 1.3] that $\overline{\varphi_{G}^{-1}}\circ\overline{\varphi_{G}}(P)=P$ and $\overline{\varphi_{G}}\circ\overline{\varphi_{G}^{-1}}(Q)=Q$ hold for all $P\in\mathcal{P}^{+}(^{n}E;G^{\ast})$ and $Q\in\mathcal{P}^{+}(^{n}F;G^{\ast})$. Consequently,
@@
In [21], E. Lacey constructed a nonatomic AM-space $X$ that is linearly isomorphic to $C\langle 1,\alpha\rangle$, where $\alpha\geq\omega$ is a countable ordinal and $\langle 1,\alpha\rangle$ denotes the order interval of ordinals between $1$ and $\alpha$ with the usual order topology. Moreover, there exists an isometric isomorphism $R\colon X^{*}\to\ell_{1}$ that is also order preserving, which means that $R\varphi\geq 0$ in $\ell_{1}$ if and only if $\varphi\geq 0$ in $X^{*}$. This yields that $R$ preserves the lattice structure, and hence $R$ is a lattice isomorphism (see [20, p. 5] or [19, Example 3.1]), proving that $X^{*}$ and $\ell_{1}$ are lattice isomorphic. Let us check that $c_{0}^{\ast}$ and $c^{\ast}$ are also lattice isomorphic to $\ell_{1}$:
@@
(ii) Let $T\colon\ell_{1}\to c^{*}$ be the linear isomorphism given by $T(b)(x)=b_{1}\displaystyle\lim_{n\to\infty}x_{n}+\sum_{j=2}^{\infty}b_{j}x_{j-1}$ for all $b=(b_{j})_{j}\in\ell_{1}$ and $x=(x_{j})_{j}\in c$. Clearly, $T\geq 0$. The inverse map $T^{-1}\colon c^{*}\to\ell_{1}$ is such that $T^{-1}(\varphi)=(\varphi(e)-\displaystyle\sum_{j=1}^{\infty}\varphi(e_{j}),\varphi(e_{1}),\varphi(e_{2}),\dots)$, where $e=(1,1,1,\dots)$. Indeed, for every $x=(x_{j})_{j}\in c$, we have that $x=(\displaystyle\lim_{n\to\infty}x_{n})\,e\,+\,\sum_{j=1}^{\infty}(x_{j}-\lim_{n\to\infty}x_{n})e_{j},$ and hence
@@
We begin recalling from [7, Example 2.3] that, under the assumption on $G^{\ast}$, $\mathcal{P}_{\mathcal{K}}^{r}(^{n}E;G^{\ast})$ and $\mathcal{P}_{\mathcal{K}}^{r}(^{n}F;G^{\ast})$ are closed sublattices of $\mathcal{P}^{r}(^{n}E;G^{\ast})$ and $\mathcal{P}^{r}(^{n}F;G^{\ast})$, respectively. Letting $\overline{\varphi_{G}}\colon\mathcal{P}^{r}(^{n}E;G^{\ast})\to\mathcal{P}^{r}(^{n}F;G^{\ast})$ be the lattice isomorphism defined in the proof of Theorem 2.5, we claim that $\overline{\varphi_{G}}(P)$ is a compact polynomial for every $P\in\mathcal{P}_{\mathcal{K}}^{+}(^{n}E;G^{\ast})$:
@@
$\lim_{j\to\infty}\overline{\varphi_{G}}(P)(y_{k_{j}})=\displaystyle\lim_{j\to\infty}\overline{P}(\varphi^{\ast}(J_{F}(y_{k_{j}})))|_{G}=g^{\ast\ast\ast}|_{G}\in G^{\ast},$
@@
The following holds by applying the same argument used above replacing [7, Example 2.3] with [7, Example 2.4].
Corollary 2.9.
@@
Given Banach lattices $E$, $F$. and $n\in\mathbb{N}$, we say that an $n$-homogeneous polynomial $P\in\mathcal{P}^{r}(^{n}E;F)$ is orthogonally additive if $P(x+y)=P(x)+P(y)$ whenever $|x|\wedge|y|=0$; or equivalently if $T_{P}$ is orthosymmetric, that is $T_{P}(x_{1},\dots,x_{m})=0$ whenever some pair $x_{i}$, $x_{j}$ of the arguments are disjoint, i.e. $|x_{i}|\wedge|x_{j}|=0$ (see [11, Lemma 4.1]). In the case where $F$ is Dedekind complete, the Banach lattice of all regular polynomials $P\colon E\to F$ for which $|P|$ is orthogonally additive is denoted by $\mathcal{P}_{o}^{r}(^{n}E;F)$ (see [11, Section 5]).
@@
Let $\varphi\colon E^{*}\to F^{*}$ be the lattice isomorphism given by the assumption. By [1, Theorem 2.20] we get that $\varphi^{*}$ is interval preserving. Since $\varphi$ is surjective, $\varphi^{*}$ is also injective. Thus, $\varphi^{*}$ is a lattice homommorphism by [30, Proposition 2.1]. Letting $\overline{\varphi_{G}}\colon\mathcal{P}^{r}(^{n}E;G^{\ast})\to\mathcal{P}^{r}(^{n}F;G^{\ast})$ be the lattice isomorphism defined in the proof of Theorem 2.5, we claim that $\overline{\varphi_{G}}(P)$ is an orthogonally additive polynomial for every $P\in\mathcal{P}_{o}^{+}(^{n}E;G^{\ast})$. Indeed, given $y_{1},\dots,y_{n}\in F$ with $|y_{i}|\wedge|y_{j}|=0$ for some $i\neq j$, we have that $|\varphi^{*}(J_{F}(y_{i}))|\wedge|\varphi^{*}(J_{F}(y_{j}))|=0$, because both $\varphi^{*}$ and $J_{F}$ preserve disjointness. Since the Arens-extension of an order bounded orthosymmetric $n$-linear map is also orthosymmetric (see [28, Corollary 5.18]), we conclude that
@@
proving that $T_{\overline{\varphi_{G}}(P)}$ is orthosymmetric. Thus, $\overline{\varphi_{G}}(P)\in\mathcal{P}_{o}^{+}(^{n}F;G^{*})$ as we claimed. Analogously, we can prove that $\overline{\varphi_{G}^{-1}}(Q)\in\mathcal{P}_{o}^{+}(^{n}E;G^{*})$ for each $Q\in\mathcal{P}_{o}^{r}(^{n}F;G^{*})$, and therefore $\overline{\varphi_{G}}\colon\mathcal{P}_{o}^{r}(^{n}E;G^{*})\to\mathcal{P}_{o}^{r}(^{n}F;G^{*})$ is a lattice isomorphism. ∎
@@ 3 Regular nuclear polynomials
Fix $n\in\mathbb{N}$ and consider the positive bilinear map $A\colon\widehat{\otimes}_{n,s,|\pi|}E^{\ast}\,\times F\to\mathcal{P}^{r}(^{n}E;F)$ given by $A(\theta_{n}(x^{*})\otimes y)(x)=(x^{*}(x))^{n}y$ for all $0\leq x^{*}\in E^{*}$, $y\in F^{+}$, $x\in E^{+}$ and extend by linearity and continuity. Thus, $A$ induces a positive linear operator $\overline{\Phi}_{n}\colon\widehat{\otimes}_{n,s,|\pi|}E^{\ast}\,\widehat{\otimes}_{|\pi|}F\to\mathcal{P}^{r}(^{n}E;F)$ such that
@@
The elements of $\text{Im }\overline{\Phi}_{n}:=\mathcal{P}_{\mathcal{N}}^{r}(^{n}E;F)$ are called regular nuclear $n$-homogeneous polynomials. As in [4], $\mathcal{P}_{\mathcal{N}}^{r}(^{n}E;F)$ is a Banach lattice endowed with the norm $\|\cdot\|_{\mathcal{N},r}$ induced by the quotient norm on $\widehat{\otimes}_{n,s,|\pi|}E^{\ast}\,\widehat{\otimes}_{|\pi|}F/\ker\overline{\Phi}$. The main objective of this section is to prove that if $G$ is a Dedekind complete Banach lattice with the lattice approximation property and both $E^{*}$ and $F^{*}$ are lattice isomorphic, then $\mathcal{P}_{\mathcal{N}}^{r}(^{n}E;G)$ and $\mathcal{P}_{\mathcal{N}}^{r}(^{n}F;G)$ are lattice isomorphic. To do this, we recall from [4] that a principal pair of a Banach lattice $E$ is an ordered pair $(G,g)$, in which, $G$ is a finite dimensional vector sublattice of $E$ and $g$ is a strong order unit of $G$, i.e. for each $y\in G$, exists $\lambda>0$ with $|y|\leq\lambda g$. A Banach lattice $E$ is said to have the lattice approximation property (LAP in short) if for every $\varepsilon>0$ and every sequence $(E_{n},e_{n})$ of principal pairs of $E$ such that $\|e_{n}\|\to 0$ there exists finite rank operator $T\colon E\to E$ such that $\||(T-id)_{E_{n}}|(e_{n})\|<\varepsilon$ for every $n\in\mathbb{N}$.
The following lemma will be necessary to achieve our objective.
@@
Let $E$ and $F$ be two Banach lattices with $F$ being Dedekind complete. For every $w\in\widehat{\otimes}_{n,s,|\pi|}E^{\ast}\,\widehat{\otimes}_{|\pi|}F$, there exist a sequence $(F_{k})_{k}$ of finite-dimensional vector sublattices of $F$ and a sequence $(w_{k})_{k}$ in $\widehat{\otimes}_{n,s,|\pi|}E^{\ast}\,\widehat{\otimes}_{|\pi|}F$ such that $\|w\|_{|\pi|}=\sum_{k=1}^{\infty}\|w_{k}\|_{|\pi|}<\infty$, $w=\sum_{k=1}^{\infty}w_{k}$, $w_{k}\in\widehat{\otimes}_{n,s,|\pi|}E^{\ast}\otimes F_{k}$ and $\|w_{k}\|_{\widehat{\otimes}_{n,s,|\pi|}E^{\ast}\,\widehat{\otimes}_{|\pi|}F_{k}}\leq 2\|w_{k}\|_{|\pi|}+2^{1-k}$ for every $k\in\mathbb{N}$. Moreover, $\|w_{k}\|_{\widehat{\otimes}_{n,s,|\pi|}E^{\ast}\,\widehat{\otimes}_{|\pi|}F_{k}}\leq\frac{6}{2^{k}}$ holds for every $k\in\mathbb{N}$ whenever $\|w\|_{|\pi|}=1$.

### v2

@@ 2 Díaz-Dineen problem for regular polynomials
is positive, symmetric (because $\overline{T_{P}}$ is symmetric by the assumption and Corollary 2.2) and satisfies $\overline{\varphi_{G}}(P)(y)=A(y,\dots,y)$. Thus $\overline{\varphi_{G}}(P)\geq 0$ for every positive $n$-homogeneous polynomial $P\colon E\to G$. Therefore, we may consider the positive linear operator $\overline{\varphi_{G}}$ from $\mathcal{P}^{r}(^{n}E;G^{\ast})$ into $\mathcal{P}^{r}(^{n}F;G^{\ast})$. Analogously, since $\varphi^{-1}\colon F^{\ast}\to E^{\ast}$ is a positive operator (see [1, Theorem 2.15]), we get that $\overline{\varphi_{G}^{-1}}\colon\mathcal{P}^{r}(^{n}F;G^{\ast})\to\mathcal{P}^{r}(^{n}E;G^{\ast})$ is also a positive operator.
Finally, since $E^{*}$ and $F^{*}$ have order continuous norm, we get by the same argument used in the proof of [17, Theorem 1.3] that $\overline{\varphi_{G}^{-1}}\circ\overline{\varphi_{G}}(P)=P$ and $\overline{\varphi_{G}}\circ\overline{\varphi_{G}^{-1}}(Q)=Q$ hold for all $P\in\mathcal{P}^{+}(^{n}E;G^{\ast})$ and $Q\in\mathcal{P}^{+}(^{n}F;G^{\ast})$. Consequently,
@@
In [28], E. Lacey constructed a nonatomic AM-space $X$ that is linearly isomorphic to $C\langle 1,\alpha\rangle$, where $\alpha\geq\omega$ is a countable ordinal and $\langle 1,\alpha\rangle$ denotes the order interval of ordinals between $1$ and $\alpha$ with the usual order topology. Moreover, there exists an isometric isomorphism $R\colon X^{*}\to\ell_{1}$ that is also order preserving, which means that $R\varphi\geq 0$ in $\ell_{1}$ if and only if $\varphi\geq 0$ in $X^{*}$. This means that $R$ preserves the lattice structure, and hence $R$ is a lattice isomorphism (see [27, p. 5] or [25, Example 3.1]), proving that $X^{*}$ and $\ell_{1}$ are lattice isomorphic. Let us check that $c_{0}^{\ast}$ and $c^{\ast}$ are also lattice isomorphic to $\ell_{1}$:
@@
(ii) Let $T\colon\ell_{1}\to c^{*}$ be the linear isomorphism given by $T(b)(x)=b_{1}\displaystyle\lim_{n\to\infty}x_{n}+\sum_{j=2}^{\infty}b_{j}x_{j-1}$ for all $b=(b_{j})_{j}\in\ell_{1}$ and $x=(x_{j})_{j}\in c$. Clearly, $T\geq 0$. The inverse map $T^{-1}\colon c^{*}\to\ell_{1}$ is given by $T^{-1}(\varphi)=(\varphi(e)-\displaystyle\sum_{j=1}^{\infty}\varphi(e_{j}),\varphi(e_{1}),\varphi(e_{2}),\dots)$, where $e=(1,1,1,\dots)$. Indeed, for every $x=(x_{j})_{j}\in c$, we have that $x=(\displaystyle\lim_{n\to\infty}x_{n})\,e\,+\,\sum_{j=1}^{\infty}(x_{j}-\lim_{n\to\infty}x_{n})e_{j},$ and hence
@@
We begin recalling from [8, Example 2.3] that, under the assumption on $G^{\ast}$, $\mathcal{P}_{\mathcal{K}}^{r}(^{n}E;G^{\ast})$ and $\mathcal{P}_{\mathcal{K}}^{r}(^{n}F;G^{\ast})$ are closed sublattices of $\mathcal{P}^{r}(^{n}E;G^{\ast})$ and $\mathcal{P}^{r}(^{n}F;G^{\ast})$, respectively. Letting $\overline{\varphi_{G}}\colon\mathcal{P}^{r}(^{n}E;G^{\ast})\to\mathcal{P}^{r}(^{n}F;G^{\ast})$ be the lattice isomorphism defined in the proof of Theorem 2.4, we claim that $\overline{\varphi_{G}}(P)$ is a compact polynomial for every $P\in\mathcal{P}_{\mathcal{K}}^{+}(^{n}E;G^{\ast})$:
@@
$\lim_{j\to\infty}\overline{\varphi_{G}}(P)(y_{k_{j}})=\displaystyle\lim_{j\to\infty}\overline{P}(\varphi^{\ast}(J_{F}(y_{k_{j}})))|_{G}=g^{\ast}\in G^{\ast},$
@@
The following holds by applying the same argument used above replacing [8, Example 2.3] with [8, Example 2.4].
Proposition 2.8.
@@
Given Banach lattices $E$, $F$, and $n\in\mathbb{N}$, we say that an $n$-homogeneous polynomial $P\in\mathcal{P}^{r}(^{n}E;F)$ is orthogonally additive if $P(x+y)=P(x)+P(y)$ whenever $|x|\wedge|y|=0$; or equivalently if $T_{P}$ is orthosymmetric, that is $T_{P}(x_{1},\dots,x_{m})=0$ whenever some pair $x_{i}$, $x_{j}$ of the arguments are disjoint, i.e. $|x_{i}|\wedge|x_{j}|=0$ (see [14, Lemma 4.1]). In the case where $F$ is Dedekind complete, the Banach lattice of all regular polynomials $P\colon E\to F$ for which $|P|$ is orthogonally additive is denoted by $\mathcal{P}_{o}^{r}(^{n}E;F)$ (see [14, Section 5]).
@@
Let $\varphi\colon E^{*}\to F^{*}$ be the lattice isomorphism given by the assumption. By [1, Theorem 2.20] we get that $\varphi^{*}$ is interval preserving. Since $\varphi$ is surjective, $\varphi^{*}$ is also injective. Thus, $\varphi^{*}$ is a lattice homomorphism by [38, Proposition 2.1]. Letting $\overline{\varphi_{G}}\colon\mathcal{P}^{r}(^{n}E;G^{\ast})\to\mathcal{P}^{r}(^{n}F;G^{\ast})$ be the lattice isomorphism defined in the proof of Theorem 2.4, we claim that $\overline{\varphi_{G}}(P)$ is an orthogonally additive polynomial for every $P\in\mathcal{P}_{o}^{+}(^{n}E;G^{\ast})$. Indeed, given $y_{1},\dots,y_{n}\in F$ with $|y_{i}|\wedge|y_{j}|=0$ for some $i\neq j$, we have that $|\varphi^{*}(J_{F}(y_{i}))|\wedge|\varphi^{*}(J_{F}(y_{j}))|=0$, because both $\varphi^{*}$ and $J_{F}$ preserve disjointness. Since the Arens-extension of an order bounded orthosymmetric $n$-linear map is also orthosymmetric (see [35, Corollary 5.18]), we conclude that
@@
proving that $T_{\overline{\varphi_{G}}(P)}$ is orthosymmetric. Thus, $\overline{\varphi_{G}}(P)\in\mathcal{P}_{o}^{+}(^{n}F;G^{*})$ as we claimed. Analogously, we can prove that $\overline{\varphi_{G}^{-1}}(Q)\in\mathcal{P}_{o}^{+}(^{n}E;G^{*})$ for each $Q\in\mathcal{P}_{o}^{+}(^{n}F;G^{*})$, and therefore $\overline{\varphi_{G}}\colon\mathcal{P}_{o}^{r}(^{n}E;G^{*})\to\mathcal{P}_{o}^{r}(^{n}F;G^{*})$ is a lattice isomorphism. ∎
@@ 3 Regular nuclear polynomials
To present a polynomial version of such class, fix $n\in\mathbb{N}$ and let $\phi_{1},\dots,\phi_{n}\in(E^{*})^{+}$ be given. If $A:E\times\overset{n}{\cdots}\times E\to\mathbb{R}$ is the $n$-linear operator defined by
$A(x_{1},\dots,x_{n})=\phi_{1}(x_{1})\cdots\phi_{n}(x_{n}),$
its symmetrization is given by
$A^{s}(x_{1},\dots,x_{n})=\frac{1}{n!}\sum_{\sigma\in S_{n}}A(x_{\sigma(1)},\dots,x_{\sigma(n)})=\frac{1}{n!}\sum_{\sigma\in S_{n}}\phi_{1}(x_{\sigma(1)})\cdots\phi_{n}(x_{\sigma(n)}).$
Since $\phi_{1},\dots,\phi_{n}\geq 0$, we get that $A^{s}\geq 0$. Thus $P_{\phi_{1},\dots,\phi_{n}}(x):=A(x,\dots,x)=\phi_{1}(x)\cdots\phi_{n}(x)$ defines a positive $n$-homogeneous polynomial on $E$, and so we can define a positive symmetric $n$-linear map
$(\phi_{1},\dots,\phi_{n})\in E^{*}\times\cdots\times E^{*}\mapsto P_{\phi_{1},\dots,\phi_{n}}\in\mathcal{P}^{r}(^{n}E).$
Its range is contained in $\mathcal{P}^{r}(^{n}E)$. For instance, taking $n=2$,
$\phi_{1}\cdot\phi_{2}=(\phi_{1}^{+}-\phi_{1}^{-})(\phi_{2}^{+}-\phi_{2}^{-})=(\phi_{1}^{+}\phi_{2}^{+}+\phi_{1}^{-}\phi_{2}^{-})-(\phi_{1}^{+}\phi_{2}^{-}+\phi_{1}^{-}\phi_{2}^{+}).$
Now, we can consider the positive bilinear map $B\colon\widehat{\otimes}_{n,s,|\pi|}E^{\ast}\times F\to\mathcal{P}^{r}(^{n}E;F)$ given by $B(\theta_{n}(x^{*})\otimes y)(x)=T(\theta_{n}(x^{*}))y=(x^{*}(x))^{n}y$ for all $0\leq x^{*}\in E^{*}$, $y\in F^{+}$, $x\in E^{+}$. Hence, $B$ induces a positive linear operator $\overline{\Phi}_{n}\colon\widehat{\otimes}_{n,s,|\pi|}E^{\ast}\,\widehat{\otimes}_{|\pi|}F\to\mathcal{P}^{r}(^{n}E;F)$ such that
@@
The elements of $\text{Im }\overline{\Phi}_{n}:=\mathcal{P}_{\mathcal{N}}^{r}(^{n}E;F)$ are called regular nuclear $n$-homogeneous polynomials. As in [4], $\mathcal{P}_{\mathcal{N}}^{r}(^{n}E;F)$ is a Banach lattice endowed with the norm $\|\cdot\|_{\mathcal{N},r}$ induced by the quotient norm on $\widehat{\otimes}_{n,s,|\pi|}E^{\ast}\,\widehat{\otimes}_{|\pi|}F/\ker\overline{\Phi}$. The main objective of this section is to prove that if $G$ is a Dedekind complete Banach lattice with the lattice approximation property, and both $E^{*}$ and $F^{*}$ have the lattice approximation property and are lattice isomorphic, then $\mathcal{P}_{\mathcal{N}}^{r}(^{n}E;G)$ and $\mathcal{P}_{\mathcal{N}}^{r}(^{n}F;G)$ are lattice isomorphic. To do this, we recall from [4] that a principal pair of a Banach lattice $E$ is an ordered pair $(G,g)$, in which, $G$ is a finite dimensional vector sublattice of $E$ and $g$ is a strong order unit of $G$, i.e. for each $y\in G$, exists $\lambda>0$ with $|y|\leq\lambda g$. A Banach lattice $E$ is said to have the lattice approximation property (LAP in short) if for every $\varepsilon>0$ and every sequence $(E_{n},e_{n})$ of principal pairs of $E$ such that $\|e_{n}\|\to 0$ there exists finite rank operator $T\colon E\to E$ such that $\||(T-id)_{E_{n}}|(e_{n})\|<\varepsilon$ for every $n\in\mathbb{N}$.
The following lemma, which can be seen as a generalization of [4, Lemma 3.5], will be necessary to achieve our objective.
@@
Let $E$ and $F$ be Dedekind complete Banach lattices and $w\in\widehat{\otimes}_{n,s,|\pi|}E\,\widehat{\otimes}_{|\pi|}\,F$. For each $0<\varepsilon\leq 1$, there exist finite-dimensional vector sublattices $E_{0}\subset E$ and $F_{0}\subset F$ and a vector $\widetilde{w}\in\otimes_{n,s}E_{0}\,\otimes F_{0}$ such that $\|w-\tilde{w}\|_{s,|\pi|}\leq\varepsilon$ and $\|\tilde{w}\|_{\otimes_{n,s}E_{0}\widehat{\otimes}_{|\pi|}F_{0}}\leq(1+\varepsilon)^{n}\|\tilde{w}\|_{s,|\pi|}+2^{n}\varepsilon$